OpenStudy (anonymous):

Which ordered pair is NOT a solution of the inequality, 3x - 2y > (there is a line that goes underneath the arrow) 12 A. (-3,-10) B. (-1,-7) C. (0,-6) D. (2,-3)

OpenStudy (anonymous):

\[3x-2y \ge 12\]

OpenStudy (anonymous):

just substiture the points in x and y and you'll get the answer. It's B

OpenStudy (anonymous):

Thats what I thought, but for some reason I thought that it was D

OpenStudy (anonymous):

3(2)-2(-3) = 6+6 which is equal to 12 so it satisfies the inequality